OpenStudy (anonymous):

Which of the following is the formula for the geometric sequence? 320, 80, 20, 5,... A. an = 320(1/4)n-1 B. an = 320(4)n-1 C. an = 4(320)n-1

OpenStudy (kinggeorge):

Do you have a guess as to which one you think is correct?

OpenStudy (across):

It would not hurt to plug in a few values into each function. :P

OpenStudy (anonymous):

b?

sam (.sam.):

If geometric, r=80/320=1/4 a=320 formula gives \[\huge T_n=ar^{n-1}\] \[\huge T_n=320(1/4)^{n-1}\]

OpenStudy (anonymous):

thankyou!!!:)